# CoBRA: A Universal Strategyproof Confirmation Protocol for Quorum-based Proof-of-Stake Blockchains

### v2

@@ 2.3 SMR game
We partition the space of strategy profiles into two sets: $S_{\text{bad}}$, consisting of strategy profiles resulting in safety violations (i.e., conflicting blocks are finalized with valid certificates) or liveness violation (i.e., permanent censorship of a transaction), and $S_{good}$ which preserve both properties.
For each validator $p$, we define $S_{\mathsf{bad}}^{p}$ as the set of individual strategies $s_{p}$ for which there exists a strategy profile $s=(s_{p},s_{-p})\in S_{\mathsf{bad}}$ where $p$ actively contributes to the violation. Conversely, let $S_{\mathsf{good}}$ be the set of strategy profiles that preserve both safety and liveness. For each validator $p$, denote by $S_{\mathsf{good}}^{p}$ the set of individual strategies $s_{p}$ such that there exists a strategy profile $s=(s_{p},s_{-p})\in S_{\mathsf{good}}$.
We say that a protocol is rationally robust if, for every rational validator $p$, any strategy $s_{p}\in S_{\mathsf{bad}}^{p}$ that contributes to a violation is weakly dominated by some strategy $s_{p}^{\prime}\in S_{\mathsf{good}}^{p}$. In other words, rational validators always have an equally good or better option than causing violations.
@@
A strategy $s_{i}\in S_{i}$ weakly dominates a strategy $s_{i}^{\prime}\in S_{i}$ if, for every possible strategy profile $s_{-i}\in S_{-i}$, the following holds:
$u_{i}(s_{i},s_{-i})\geq u_{i}(s^{\prime}_{i},s_{-i}),$
and there exists at least one $s_{-i}\in S_{-i}$ such that:
$u_{i}(s_{i},s_{-i})>u_{i}(s^{\prime}_{i},s_{-i}).$
@@
We say that an SMR game is rationally robust if for every rational validator $p$, and every strategy $s_{p}\in S_{\mathsf{bad}}^{p}$, there exists a strategy $s_{p}^{\prime}\in S_{\mathsf{good}}^{p}$ that weakly dominates $s_{p}$.
Rational robustness relaxes traditional incentive compatibility, which requires that following the protocol is always the utility-maximizing strategy. Such a requirement is too strong in adversarial settings, where Byzantine validators may accumulate participation rewards and use them to bribe rational ones. Instead, rational robustness focuses only on the minimal conditions to guarantee safety and liveness.
3 Impossibility results
In the following, we identify settings in which a $q$-committable $(n,f)$–resilient PoS blockchain cannot be transformed into an $(n,k,f)$–resilient PoS blockchain. Additionally, we determine when $q$-committable SMR protocols fail to guarantee finality before the synchrony bound $\Delta^{*}$ elapses. We provide a summary of our results along with proof sketches and defer the complete proofs to Appendix C.
@@ 3 Impossibility results
In Theorem 3, we show that designing an $(n,k,f)$-resilient $q$-commitable SMR protocol is impossible for $f\geq\lceil\frac{h}{2}\rceil$, where $h$ is the number of correct validators. Notice that this result depends only on the proportion of Byzantine and correct validators. This is because to guarantee liveness, the size of the quorum $q$ is at most $n-f$. For $f\geq\lceil\frac{h}{2}\rceil$, we cannot guarantee that two quorums intersect in at least one correct validator. Rational validators, in turn, will collude with Byzantine validators to form quorums of conflicting blocks to maximize their monetary gains.
To establish this result, we extend the proof of Budish et al. [12, Theorem 4.2] within our model. The core issue is that rational validators can collude with Byzantine to generate valid quorum certificates while avoiding penalties. This is possible because, for any predefined withdrawal period that allows validators to reclaim their stake, correct validators cannot reliably distinguish between an honest execution with network-induced delays and malicious behavior. Hence, rational validators may engage in equivocation without facing slashing, undermining the protocol’s security.
@@
theoremPartialSynchrony Assuming silent clients, a static setting, and a partially synchronous network, no $q$-commitable SMR protocol can be $(n,k,f)$–resilient for any $q$ when $f\geq\lceil\frac{h}{2}\rceil$, where $h=n-k-f$.
Proof sketch of Theorem 3
Proof.
The adversary can partition correct validators in two distinct subsets $H_{1}$, $H_{2}$ and perform the double-spending attack described in Lemma 8.This attack can be beneficial for rational validators since for any value of stake $D$ they lock before the run, there is an input of conflicting transactions $T_{1}$ and $T_{2}$ where rational validators transfer all their liquid coins to distinct subset of clients $C_{1}$ and $C_{2}$ respectively. Notably, misbehaving validators can withdraw their stake without even being slashed.
It follows directly from Theorem 3 that no $q$-commitable protocol can be $(n,f)$–resilient when $f=\lceil h/2\rceil$, implying resilience is impossible if Byzantine validators exceed one-third of the network ($f\geq n/3$). Furthermore, even if a $q$-commitable protocol is resilient for $f=\lfloor h/2\rfloor$, replacing just one correct validator with a rational one breaks resilience if the total number of Byzantine and rational validators reaches at least one third the network ($f+k\geq n/3$), as shown in Corollary 3.1. This result highlights the fragility of quorum-based protocols under rational behavior, as even a small shift from correct to rational participants significantly impacts security.
@@
Consider any $q$-commitable protocol $\Pi$ that is $(n,f)$–resilient for $f=\lfloor\frac{h}{2}\rfloor$ where $h=n-f$ is the number of correct validators. For any $k\geq 1$, $\Pi$ is not $(n,k,f)$–resilient in our model in the partial synchronous setting.
Corollary 3.1 implies that no quorum-based SMR protocol $\Pi$ can achieve $(3f+1,k,f)$–resilience for $k\geq 1$ in the partial synchronous model. That means that quorum-based SMR protocols are not resilient for $f+k\geq n/3$ under the model described in Theorem 3.
@@
theoremSynchrony Assuming silent clients, a static setting, and a synchronous network, no $q$-commitable SMR protocol can be $(n,k,f)$-resilient, when $f\geq h$, where $h=n-k-f$.
Proof sketch of Theorem 3
Proof.
When $f+k\geq q$, Byzantine and rational validators can successfully construct valid certificates for an arbitrary number of conflicting blocks and provide them to clients. For any stake $D$ that rational validators have locked before the run, there is a configuration in which each rational validator can double-spend more than $D$ coins. That is because the number of clients is arbitrary and unknown before the run. Even if the conflicting blocks are detected and the misbehaving validators get slashed, the attack is still beneficial for rational validators. ∎
By setting $k=0$ in Theorem 3, we conclude that no $q$-commitable protocol can be $(n,f)$–resilient when $f=h$. Furthermore, Theorem 3 establishes that even if a $q$-commitable protocol remains $(n,f)$–resilient for $f=h-1$, we cannot replace a single correct validator to a rational one, implying that resilience breaks down when $f+k\geq 2n/3$. This result is formalized in Corollary 3.2.
@@
Consider any $q$-commitable protocol $\Pi$ that is $(n,f)$–resilient for $f=h-1$ where $h=n-f$ is the number of correct validators. For any $k\geq 1$, $\Pi$ is not $(n,k,f)$–resilient under our model in the synchronous setting.
In other words, no quorum-based SMR protocol can achieve $(n,f+1,f)$–resilience, even in the synchronous model. Therefore, quorum-based SMR protocols are not resilient for $f+k\geq 2n/3$.
@@
Theorem 3 establishes that if participation rewards are used to incentivize rational validators, no protocol can ensure finality during a $\Delta^{*}$-bounded window when $f\geq\lceil\frac{h}{2}\rceil$. That is because Byzantine and rational validators can still produce valid quorum certificates for conflicting blocks within $\Delta^{*}$, with each rational validator double-spending an amount exceeding their stake. This attack remains profitable for rational validators, even if they get slashed and forfeit their stake.
@@
theoremresponsiveness Let us assume participation rewards are distributed to validators, silent clients, a static PoS setting, and a synchronous network. No SMR protocol can be both (i) $(n,f)$–resilient $q$-commitable with finality latency $t$, and (ii) $(n,k,f)$–resilient with finality latency $t^{\prime}<t+\Delta^{*}$, when $f\geq\lceil\frac{h}{2}\rceil$, for $h=n-k-f$.
Proof sketch of Theorem 3
Proof.
The adversary can perform the double-spending attack illustrated in Lemma 8, within the last $\Delta^{*}$. This attack can be beneficial for rational validators since for any value of stake $D$ they lock before the run, there is an input of conflicting transactions $T_{1}$ and $T_{2}$ where rational validators transfer $D+\epsilon,\epsilon>0$ coins to clients in $C_{1}$ and $C_{2}$ respectively, and therefore they can double-spend more than $D$ coins. Even if they are slashed and forfeit their stake, the attack is still beneficial.
Now, let us fix any value of stake $D$. Similar to Theorem 3, a safety violation remains possible within $\Delta^{*}$. That is because both cases assume the same resilience against Byzantine adversaries and operate under similar network conditions during a network partition in $\Delta^{*}$.
We only need to argue that we can construct a configuration where validators can spend more than $D$ coins in $T_{1}$ and $T_{2}$. To show that, we argue that validators can spend their funds multiple times. This is feasible because they can reclaim their initial funds through participation rewards, allowing them to spend them again, and repeat that sufficiently many times during the period $\Delta^{*}$.
@@
According to Theorem 3, if we run an $(n,f)$–resilient quorum-based SMR protocol $\Pi$ with $f\geq\lceil\frac{h}{2}\rceil$ validators, it is impossible to finalize every transaction before $\Delta^{*}$ elapses.
4 CoBRA
In this section, we outline how to transform any $(n,f)$-resilient $(2f+1)$-commitable SMR protocol $\Pi$ into a $(n,k,f^{*})$-resilient proof-of-stake SMR protocol for $n=3f+1$. First, we consider $f^{*}\leq f$ Byzantine validators, preserving the security guarantees of the underlying SMR protocol while allowing the model to accommodate more rational nodes. To accommodate at least one rational validator ($k\geq 1$), we assume a synchronous network model with message delays bounded by $\Delta^{*}$, as required by Theorem 3. Under these conditions, we provide an $(n,k,f^{*})$-resilient protocol with $k+f^{*}\leq 2f$ and $f^{*}\leq f$, which is optimal as shown in Theorem 3.
4.1 CoBRA Overview
Our protocol builds on an underlying SMR protocol $\Pi$ and ensures that validators finalize transactions securely. The key challenge is that rational validators may attempt to maximize their gains by colluding with Byzantine validators to double-spend or censor transactions. Furthermore, as proven in Theorem 3, no protocol can ensure finality before the maximum message delay elapses. To address these challenges, our protocol introduces the following components:
  • Safety. A finality gadget that defines confirmation rules for validators and clients, and detects equivocations; together with a slashing mechanism that penalizes validators upon evidence of equivocation. These components ensure that rational validators have no incentive to fork, as such behavior yields no reward.

### v3

@@ 2.3. SMR Game and Coalition-Compliance
We partition the space of strategy profiles into two sets: $S_{\text{bad}}$, consisting of strategy profiles resulting in safety violations (i.e., conflicting blocks are finalized with valid certificates) or liveness violations (i.e., permanent censorship of a transaction), and $S_{\text{good}}$ which preserve both properties.
For each validator $p$, we define $S_{\mathsf{bad}}^{p}$ as the set of individual strategies $s_{p}$ for which there exists a strategy profile $s=(s_{p},s_{-p})\in S_{\mathsf{bad}}$ where $p$ actively contributes to the violation. Conversely, for each validator $p$, we denote by $S_{\mathsf{good}}^{p}$ the set of individual strategies $s_{p}$ such that there exists a strategy profile $s=(s_{p},s_{-p})\in S_{\mathsf{good}}$.
We say that a protocol is coalition-compliant if no rational validator can profit from contributing to safety or liveness violations, even under full collusion with Byzantine validators. In other words, rational validators always have a protocol-compliant strategy that is equally good or better than any deviation causing violations.
@@
A strategy $s_{i}\in S_{i}$ weakly dominates a strategy $s_{i}^{\prime}\in S_{i}$ if, for every possible strategy profile $s_{-i}\in S_{-i}$, it holds that $u_{i}(s_{i},s_{-i})\geq u_{i}(s^{\prime}_{i},s_{-i}),$ and there exists at least one $s_{-i}\in S_{-i}$ such that: $u_{i}(s_{i},s_{-i})>u_{i}(s^{\prime}_{i},s_{-i}).$
@@
We say that an SMR game is $(n,k,f)$-coalition-compliance if for any rational validator $p$, and every strategy $s_{p}\in S_{\mathsf{bad}}^{p}$, there exists a strategy $s_{p}^{\prime}\in S_{\mathsf{good}}^{p}$ that weakly dominates $s_{p}$.
Coalition-compliance is a stronger notion than equilibrium, as it accounts for arbitrary collusion among subsets of rational and Byzantine validators. However, coalition-compliance relaxes traditional incentive compatibility, which requires that following the protocol is always the utility-maximizing strategy. Such a requirement is too strong in adversarial settings, where Byzantine validators may accumulate participation rewards and use them to bribe rational ones. Instead, rational coalition-compliance focuses only on the minimal conditions to guarantee safety and liveness.
3. Impossibility results
In the following, we identify settings in which a $q$-committable (n,f)(n,f)–resilient PoS blockchain cannot be transformed into an $(n,k,f^{*})$–resilient PoS blockchain for $f^{*}\leq f$. Additionally, we determine when $q$-committable SMR protocols fail to guarantee finality before the synchrony bound $\Delta^{*}$ elapses.
We provide a summary of our results and defer the proof sketches along with the complete proofs to Appendix C. We also illustrate the main results in Figure 1. We emphasize that our impossibility results apply to the setting of silent clients, where clients decide whether to accept a block based solely on an accompanying certificate, without engaging in additional communication rounds with other participants. Finally, note that (n,k,f)(n,k,f)–coalition compliance is a deterministic guarantee which must hold in every execution of the protocol; thus, a single attack suffices to prove impossibility. We discuss probabilistic guarantees in Section 8.
@@ 3. Impossibility results
Impossibilities. (A): for $f^{*}>\lceil h/2\rceil$ (Theorem 3), (B): for $f^{*}+k>f$ (Theorem 3).
(C) Any $(n,f)$-resilient SMR protocol is also $(n,f^{*},k)$-resilient for any $f^{*}+k\leq f$.
(D) Enlarging the quorum to $q=n-f^{*}$ allows resilience for all $k\leq n-1$.
Synchrony
(A) Impossibility when $f^{*}+k\geq 2f+1$ (Theorem 3.1).
(B) Possibility of SMR only by limiting transaction volume (CoBRA, Theorem 3).
Possibilities. (C)–(D) As in the partial synchrony setting.
Figure 1. Design space of $2f+1$-committable SMR protocols.
Partial synchrony
In Theorem 3, we show that designing an $(n,k,f)$-resilient $q$-commitable SMR protocol is impossible for $f\geq\lceil\frac{h}{2}\rceil$, where $h$ is the number of correct validators. Notice that this result depends only on the proportion of Byzantine and correct validators. This is because to guarantee liveness, the size of the quorum $q$ is at most nfn-f. For $f\geq\lceil\frac{h}{2}\rceil$, we cannot guarantee that two quorums intersect in at least one correct validator. Rational validators, in turn, will collude with Byzantine validators to form quorums of conflicting blocks to maximize their monetary gains.
To establish this result, we extend the proof of Budish et al. (Budish et al., 2024a, Theorem 4.2) within our model. The core issue is that rational validators can collude with Byzantine to generate valid quorum certificates while avoiding penalties. This is possible because, for any predefined withdrawal period that allows validators to reclaim their stake, correct validators cannot reliably distinguish between an honest execution with network-induced delays and malicious behavior. Hence, rational validators may engage in equivocation without facing slashing, undermining the protocol’s security.
@@
theoremPartialSynchrony Assuming silent clients, a static setting, and a partially synchronous network, no $q$-commitable SMR protocol can be (n,k,f)(n,k,f)–resilient for any $q$ when $f\geq\lceil\frac{h}{2}\rceil$, where h=nkfh=n-k-f.
It follows directly from Theorem 3 that no $q$-commitable protocol can be (n,f)(n,f)–resilient when $f=\lceil h/2\rceil$, which agrees with the well-known result that resilience is impossible once Byzantine validators exceed one third of the network ($f\geq n/3$). Furthermore, the additional insight provided by our result is that even if a $q$-commitable protocol is resilient for $f=\lfloor h/2\rfloor$, replacing just one correct validator with a rational one breaks resilience if the total number of Byzantine and rational validators reaches at least one third the network ($f+k\geq n/3$), as shown in Corollary 3.1. This result highlights the fragility of quorum-based protocols under rational behavior, as even a small shift from correct to rational participants significantly impacts security.
@@
Consider any $q$-commitable protocol $\Pi$ that is (n,f)(n,f)–resilient for $f=\lfloor\frac{h}{2}\rfloor$ where h=nfh=n-f is the number of correct validators. For any $k\geq 1$, $\Pi$ is not (n,k,f)(n,k,f)–resilient in our model in the partial synchronous setting.
Corollary 3.1 implies that no quorum-based SMR protocol $\Pi$ can achieve (3f+1,k,f)(3f+1,k,f)–resilience for $k\geq 1$ in the partial synchronous model. That means that quorum-based SMR protocols are not resilient for $f+k\geq n/3$ under the model described in Theorem 3.
Fewer Byzantine.
Theorem 3 characterizes the maximum Byzantine resilience of $(n,k,f)$-resilient SMR protocols, independently of the quorum size $q$. A natural question is what happens below this maximum Byzantine threshold. Can an SMR protocol remain secure when $f^{*}<\lfloor h/2\rfloor$ and the combined number of Byzantine and rational validators exceeds one third of the network, i.e., $f^{*}+k>n/3$?
When $f=\lfloor h/2\rfloor$, even a single rational validator can collude with Byzantine validators to equivocate without being detected by other rational validators. However, for $f^{*}<\lfloor h/2\rfloor$, more than one rational validators must collude to equivocate. This opens a potential design space in which rational validators can be incentivized to expose misbehavior. One such approach is a baiting strategy, leveraged by TRAP (Ranchal-Pedrosa and Gramoli, 2022) to achieve consensus on a single value, where rational validators are rewarded for reporting equivocation. Theorem 3 characterizes the conditions under which, even in the presence of baiting strategies, solving SMR remains impossible.
{restatable}
theorempartialSync2 Let us assume participation rewards are distributed to validators, silent clients, a static PoS setting, and a partially synchronous network. Without any bounds i) on the liquid funds held by rational validators or malicious clients, ii) or on the participation rewards, the following holds: no $q$-commitable SMR protocol can be (n,k,f)(n,k,f)–resilient for any $q$ when $q\leq f+k+\lfloor\frac{h}{2}\rfloor$, where h=nkfh=n-k-f.
Theorem 3 implies that no quorum-based SMR protocol can achieve $(3f+1,k,f^{*})$–resiliency with $k=f-f^{*}+1$ in partial synchrony without enlarging the quorum size, as formalized below.
Corollary 3.2.
Consider any $q$-commitable protocol $\Pi$ that is (n,f)(n,f)–resilient for $f=\lfloor\tfrac{h}{2}\rfloor$, where h=nfh=n-f is the number of correct validators. For any $f^{*}<f$ and $k=f-f^{*}+1$, the protocol $\Pi$ is not $(n,k,f^{*})$–resilient in our model under partial synchrony, without enlarging the quorum size.
Thus, 2f+12f+1–commitable quorum-based SMR protocols with a quorum of size $q=2f+1$ are not resilient for any $f^{*}+k\geq n/3$ under the assumptions of Theorem 3.
Closing the gaps
Corollary 3.2 establishes that, in partial synchrony, reducing the actual number of Byzantine validators does not improve resilience unless we increase the quorum size. Following the flexible BFT approach (Malkhi et al., 2019), in this setting, we can enlarge the quorum size used by clients from the classical q=nfq=n-f, to a higher quorum $q^{\prime}=n-f^{*}$, where $f^{*}$ is the actual number of Byzantine validators present. This choice ensures that $q^{\prime}>f^{*}+k+\lfloor h/2\rfloor$, so that any two quorums intersect in at least one honest validator and equivocation becomes impossible. Liveness is guaranteed by the incentives of rational validators, who are motivated to follow the protocol in order to obtain participation rewards.
@@
theoremSynchrony Assuming silent clients, a static setting, and a synchronous network, no $q$-commitable SMR protocol can be $(n,k,f)$-resilient, when $f\geq h$, where h=nkfh=n-k-f.
By setting $k=0$ in Theorem 3, we conclude that no $q$-commitable protocol can be (n,f)(n,f)–resilient when $f=h$. Furthermore, Theorem 3 establishes that even if a $q$-commitable protocol remains (n,f)(n,f)–resilient for f=h1f=h-1, we cannot replace a single correct validator to a rational one, implying that resilience breaks down when $f+k\geq 2n/3$. This result is formalized in Corollary 3.3.
@@
Consider any $q$-commitable protocol $\Pi$ that is (n,f)(n,f)–resilient for f=h1f=h-1 where h=nfh=n-f is the number of correct validators. For any $k\geq 1$, $\Pi$ is not (n,k,f)(n,k,f)–resilient under our model in the synchronous setting.
In other words, no quorum-based SMR protocol can achieve (n,f+1,f)(n,f+1,f)–resilience, even in the synchronous model. Therefore, quorum-based SMR protocols are not resilient for $f+k\geq 2n/3$.
Beyond $n/3$ Byzantine.
Some SMR protocols in synchrony can tolerate $n/3\leq f^{*}<n/2$ Byzantine validators by assuming a known bound on message delays (Abraham et al., 2020). However, the guarantees of these protocols critically depend on the network respecting this synchronous bound. In contrast, $q$-committable protocols are designed to provide security under partial synchrony, independently of any synchrony bound.
Our focus is to maintain $(n,f)$-resilience even when the network is only partially synchronous. If the synchronous bound holds, the protocol can provide additional guarantees, such as increased fault tolerance, faster finality, or improved recovery guarantees. For this reason, we do not depict this family of solutions in Figure 1.
@@
Theorem 3 establishes that without bounds on the liquid coins or the participation rewards, no protocol can ensure finality during a $\Delta^{*}$-bounded window when $f\geq\lceil\frac{h}{2}\rceil$ or or $q\leq f+k+\lceil\frac{h}{2}\rceil$. That is because Byzantine and rational validators can still produce valid quorum certificates for conflicting blocks within $\Delta^{*}$, with each rational validator double-spending an amount exceeding their stake. This attack remains profitable for rational validators, even if they get slashed and forfeit their stake.
@@
theoremresponsiveness Let us assume participation rewards are distributed to validators, silent clients, a static PoS setting, and a synchronous network. Without any bounds i) on the liquid funds held by rational validators or malicious clients, or ii) on the participation rewards, the following holds: No SMR protocol can be both (i) (n,f)(n,f)–resilient $q$-commitable with finality latency $t$, and (ii) (n,k,f)(n,k,f)–resilient with finality latency $t^{\prime}<t+\Delta^{*}$, when $f\geq\lceil\frac{h}{2}\rceil$ or $q\leq f+k+\lfloor\frac{h}{2}\rfloor$, for h=nkfh=n-k-f, for any execution.
@@
According to Theorem 3, if we run an (n,f)(n,f)–resilient quorum-based SMR protocol $\Pi$ with $f\geq\lceil\frac{h}{2}\rceil$ validators or $q\leq f+k+\lfloor\frac{h}{2}\rfloor$, it is impossible to finalize every transaction before $\Delta^{*}$ elapses. That means that we must limit the transaction volume finalized per any $\Delta^{*}$ time window.
4. CoBRA: Finalization Rule
In this section, we present the CoBRA finalization protocol, a transformation that converts any $(n,f)$-resilient $(2f+1)$-committable SMR protocol $\Pi$ into an $(n,k,f^{*})$-resilient proof-of-stake SMR protocol for $n=3f+1$.
Throughout this section, we assume a synchronous network with message delay bounded by $\Delta^{*}$, as required by Theorem 3. We consider executions in which the number of Byzantine and rational validators satisfy $f^{*}\leq f$ and $k\leq 2f-f^{*}$, respectively. Our protocol achieves the optimal resilience bounds established in Theorem 3. Recovery from executions exceeding this Byzantine threshold is deferred to Section 5.
Overview.
As established in Section 3, in the presence of rational validators and silent clients, quorum-based SMR protocols are vulnerable to two major attack vectors: (i) safety attacks, where rational validators collude with Byzantine ones to produce conflicting blocks with valid quorum certificates within the synchrony window $\Delta^{*}$, enabling double-spending attacks against clients; and (ii) censorship attacks, where the adversary uses accumulated participation rewards to bribe rational validators into withholding votes for blocks containing targeted transactions, thereby preventing their finalization.
To address these challenges, our protocol introduces the following components:
  • A finality gadget that ensures safety by enforcing stake-bounded finalization: within any $\Delta^{*}$-bounded window, the maximum transaction volume that can be finalized is limited by the total stake at risk; making equivocations unprofitable once slashing penalties are applied. This ensures that rational validators have no incentive to fork.